### v1

# IMUFace: Real-Time, Low-Power, Continuous 3D Facial Reconstruction Through Earphones

Abstract
The potential of facial expression reconstruction technology is significant, with applications in various fields such as human-computer interaction, affective computing, and virtual reality. Recent studies have proposed using ear-worn devices for facial expression reconstruction to address the environmental limitations and privacy concerns associated with traditional camera-based methods. However, these approaches still require improvements in terms of aesthetics and power consumption. This paper introduces a system called IMUFace. It uses inertial measurement units (IMUs) embedded in wireless earphones to detect subtle ear movements caused by facial muscle activities, allowing for covert and low-power facial reconstruction. A user study involving 12 participants was conducted, and a deep learning model named IMUTwinTrans was proposed. The results show that IMUFace can accurately predict users’ facial landmarks with a precision of 2.21 mm, using only five minutes of training data. The predicted landmarks can be utilized to reconstruct a three-dimensional facial model. IMUFace operates at a sampling rate of 30 Hz with a relatively low power consumption of 58 mW. The findings presented in this study demonstrate the real-world applicability of IMUFace and highlight potential directions for further research to facilitate its practical adoption.
Index Terms:
3D facial reconstruction, wearable sensing, mobile computing, low-power
I Introduction
Figure 1: Demonstration of continuous 3D facial reconstruction of IMUFace.
Facial expressions are essential in interpersonal communication because they contain various information, including emotional states [1, 2] and health conditions [3, 4]. Consequently, facial expression reconstruction is widely applied in areas requiring facial analysis. It has become a fundamental technology in fields such as human-computer interaction (HCI), virtual reality (VR), affective computing, and healthcare. For example, facial expression reconstruction can enhance users’ real-time perception of facial expressions and emotions in VR, providing a more immersive experience [5]. In metaverse applications, it can rapidly establish understanding and connections within virtual spaces [6, 7]. Additionally, it can enable silent speech interfaces, facilitating seamless human-computer interaction [8].
Early facial expression reconstruction methods typically relied on manual annotation and geometric models. These traditional methods had low automation and were complex to implement, thus facing significant limitations. With advancements in computer vision and machine learning, automated facial expression reconstruction advances the field. Traditional vision-based methods [9, 10, 11] can accurately locate facial landmarks and generate high-quality facial animations. However, the use of direct camera angles and environmental factors imposes constraints on their efficacy. In order to ensure accuracy, cameras must often face users directly, which limits the applicability of these methods in situations involving varying lighting, resolution, or occlusion conditions. Additionally, the direct capture of users’ faces may give rise to privacy concerns, particularly in contexts where 3D head movements and dynamic environments are involved, as such visual approaches are less suitable in such scenarios.
These limitations have driven a shift toward non-visual facial expression reconstruction methods. Researchers have proposed combining wearable devices and multimodal sensors to enable more flexible and privacy-friendly solutions, such as necklaces [12], hats [13], glasses [14, 15, 16] and ear-mounted devices [17, 18, 19, 20, 21]. Earlier works, like C-Face [22], used miniature ear-mounted cameras to capture facial expressions, offering flexible deployment options. Nonetheless, privacy concerns persist due to camera mobility and proximity to others. Subsequently, BioFace-3D [19] and EarIO [20] introduced non-visual 3D facial reconstruction through ear-mounted devices. However, their designs may cause discomfort by protruding behind the head or slightly touching the face. Later, EARFace [21] embedded acoustic sensors within earplug designs, yielding higher usability scores. However, high power consumption limits acoustic sensors’ viability for prolonged use. Therefore, developing a portable, unobtrusive, low-power system for real-time, continuous facial expression reconstruction remains challenging.
To address these challenges, we present IMUFace, an earphone platform equipped with two IMUs for precise facial reconstruction. These IMUs are highly sensitive, allowing them to capture subtle ear movements caused by facial muscle activity quickly and accurately.
We developed a deep learning model called IMUTwinTrans to achieve 3D facial reconstruction beyond the limitations of cameras. This model, based on the ConvTransformer architecture, is capable of simultaneously leveraging the temporal and frequency features of IMU signals to establish the correspondence between IMU signals and facial landmarks. Additionally, the model adopts a lightweight design to meet the requirements for real-time performance. During testing, the well-trained model can localize facial landmarks directly from IMU signals without visual input. The predicted facial landmarks are then filtered and fitted into a generic 3D head model to generate continuous 3D facial animation. Fig. 1 shows examples of 3D facial reconstruction using the proposed IMUFace system, demonstrating its ability to monitor various parts of the face across different facial motions.
The main contributions of this study are as follows:
  1. 1.
    We presented the first attempt to use IMU sensor signals on wireless earphones for facial landmark tracking to achieve user-friendly, continuous 3D facial expression reconstruction.
  2. 2.
    We designed a lightweight deep learning model, IMUTwinTrans, based on the ConvTransformer architecture, to achieve 3D facial reconstruction by leveraging both temporal and frequency features of IMU signal while meeting real-time performance requirements.
  3. 3.
    We optimized the power consumption of the system, enabling it to operate at a sampling rate of 30 Hz with a power consumption of only 58 mW.
II Related Work
In non-wearable 3D face reconstruction, various algorithms have been developed to address challenges posed by unconstrained photo collections and achieve high-fidelity reconstructions. These methods can be categorized based on their core algorithmic approaches: morphable models, photometric stereo techniques, and deep learning-based solutions.
Morphable models (3DMMs) have been a cornerstone in 3D face reconstruction due to their ability to represent facial geometry and texture in a compact form. Roth et al. [23] introduced an adaptive method for reconstructing 3D face surfaces from unconstrained photo collections using a 3DMM. They employed a coarse-to-fine fitting strategy that allowed the model to adapt even to low-quality image sets, thereby expanding its applicability. This approach improved the quality of reconstructions and handled variations in pose, expression, and illumination more effectively. Roth et al. further advanced this work in [24], where they proposed an algorithm capable of generating accurate 3D face models from diverse 2D images without metadata about cameras or timing. By leveraging a standard field-based Laplace editing technique, they ensured robustness across varied imaging conditions.
Photometric stereo is another key technique used in 3D face reconstruction, especially when dealing with extensive photo collections. Roth et al. [24] utilized this method to capture detailed surface normals from multiple viewpoints, enabling the construction of accurate 3D models. The integration of photometric stereo with landmark constraints provided a powerful means to drive surface reconstruction while maintaining fidelity to the input images. Moreover, Xie et al. [25] explored photometric stereo in the context of smartphone-based reconstruction, demonstrating how calibrated screen lighting can be used to recover fine-grained textures and micro facial features. Their method addresses practical limitations such as hand jitter through the use of face landmarks for alignment and outlier normal elimination techniques.
Deep learning has revolutionized 3D face reconstruction by enabling high-fidelity outputs from single images. Gecer et al. [26] introduced GANFIT, which combines Generative Adversarial Networks (GANs) and Deep Convolutional Neural Networks (DCNNs) to reconstruct facial texture and shape. GANFIT trains a powerful generator of facial texture in UV space and optimizes the parameters under the supervision of pretrained deep identity features. This approach achieves photorealistic and identity-preserving reconstructions with unprecedented detail. Built upon GANFIT, Gecer et al. [27] developed Fast-GANFIT. a faster and more robust version introducing a self-supervised regression-based initialization scheme to expedite the fitting process and enhance stability.
These studies collectively illustrate the evolution of non-wearable 3D face reconstruction algorithms, each offering unique strengths. The versatility of morphable models, the precision of photometric stereo techniques, and the fidelity of deep learning-based solutions all contribute to this advancement. This variety in methodologies highlights the continuous innovation and adaptability present in the field.
II-A 3D Face Reconstruction with Different Models
The 3D Morphable Model (3DMM) [28] serves as the foundational basis for 3D face reconstruction. It represents facial geometry through a linear combination of orthogonal basis functions, with a fixed number of points to depict a face. However, traditional 3DMM approaches require optimizing to solve for relevant coefficients, which is often complex and computationally intensive. To address these challenges, increasing attention has been directed towards leveraging deep learning techniques to directly regress these coefficients, enabling more efficient and precise 3D face reconstruction.
In the realm of deep learning, CNNs have been extensively applied to 3D face reconstruction. For instance, the paper Richardson et al. [29]presents an end-to-end CNN framework that reconstructs detailed facial geometry in a coarse-to-fine manner. Similarly, Tian et al. [30] introduce a CNN-based method for reconstructing 3D faces from two facial images, demonstrating superior performance across various datasets.
Beyond CNNs, Encoder-Decoder architectures have also been employed in 3D face reconstruction. Ji et al. [31] proposes a multi-task encoder-decoder framework based on a Siamese network to learn view-consistent shapes and textures from a single image. Meanwhile, Li et al. [32] enhanced reconstruction accuracy by using a dual-pathway encoder-decoder network to distinguish between different attributes.
Recently, Transformer models have made significant strides in 3D face reconstruction. Chen et al. [33] utilizes a conditional Generative Adversarial Network (cGAN) and a novel mesh transformer for cross-domain face synthesis and reconstruction. It introduces a reprojection consistency loss for self-supervised learning. Another study, Yaermaimaiti et al. [34] integrates ResNet and Transformer modules, leveraging self-attention mechanisms to improve parameter prediction accuracy and significantly enhance the reconstruction results.
II-B Non-wearable 3D Face Reconstruction
The research field of 3D face reconstruction employs various data sources generated by different devices to improve reconstruction quality and applicability. Traditional 3D face reconstruction methods primarily rely on external imaging devices, such as RGB cameras, utilizing extensive photo collections for reconstruction. For instance, Roth et al. [23] present a method capable of reconstructing a 3D face model of an individual from photo collections captured under varied poses and lighting conditions while also extracting albedo information. This approach achieves superior experimental results by fitting a 3D Morphable Model to form a personalized template and developing a novel photometric stereo formulation. Similarly, Roth et al. [24] introduce an algorithm that generates an accurate 3D face model from unconstrained collections of 2D images. This algorithm employs a photometric stereo method along with a novel normal field-based Laplace editing technique, leveraging images captured in diverse poses, and demonstrates both qualitative and quantitative effectiveness.
II-C Wearable 3D Face Reconstruction
In recent years, researchers have begun to explore 3D face reconstruction using wearable miniature cameras, aiming to facilitate face reconstruction in real-life scenarios while minimizing user interference. Chen et al. [22] propose an ear-mounted wearable device that captures subtle changes in facial contours using two miniature cameras and utilizes a deep learning model to continuously output facial feature points, enabling continuous facial expression reconstruction. Additionally, Aoki et al. [35] introduces a glasses-shaped device that captures the user’s facial expressions through two miniature cameras and three mirrors, allowing for the recording and recognition of facial expressions without obstructing daily activities. The system demonstrated high accuracy in emotion recognition, underscoring its applicability to facial image indexing. These studies highlight the expanding potential of wearable devices in 3D face reconstruction, opening new possibilities for practical applications.
The transition from visual methods to multi-modal approaches has led to innovative techniques for monitoring facial expressions and movements using sensors like proximity sensors, EMG, pressure sensors, EOG, and audio signals. Notably, smart glasses and earphones utilizing IMU and acoustic sensors have emerged as effective tools for 3D reconstruction. Wu et al. [19] uses a single-earpiece biosensing system to track facial landmarks and render 3D animations from biosignals without visual input, achieving accuracy comparable to camera-based solutions. Li et al. [16] employ IMUs on smart glasses to estimate action unit intensities for real-time facial reconstruction with low power consumption. Zhang et al. [21] leverages in-ear acoustic sensors to track facial landmarks by detecting changes in the ear canal shape, providing robust and privacy-preserving 3D facial reconstruction. These advancements demonstrate the potential of multi-modal sensors in offering accurate, unobtrusive, and privacy-sensitive facial tracking suitable for diverse applications.
III Background
In this section, we introduce the background information on the following aspects: (1) facial motion and the muscles involved; (2) principle of operation of IMUFace.
III-A Facial Muscles and Expressions
Facial expressions are a crucial element of human communication, providing insight into emotions and intentions that others can perceive and interpret. The human face, equipped with 43 distinct muscles, is capable of producing more than 10,000 different expressions. These muscles, known as mimetic muscles, are organized into three major categories: orbital, nasal, and oral. The orbital muscles control eyelid movements, the nasal group manages movements related to the nose, and the oral muscles facilitate motions of the lips and mouth. Through the interaction of these muscle groups, the human face can depict a wide array of expressions, such as happiness, sadness, surprise, and fear.
The complexity of facial expressions arises from the coordinated activity of these muscle groups, which pull on the facial skin to create the visual cues associated with emotions. For instance, the frontalis muscle raises the eyebrows in surprise, while the temporalis and lateral pterygoid muscles contribute to jaw movements in expressions of amazement or joy.
Figure 2: Overview of the IMUFace system.
III-B Sensing Principle
The relationship between facial expressions and the movement of facial muscles provides a foundation for innovative sensing technologies that capture these expressions indirectly. When facial muscles contract, they alter visible expressions and induce subtle changes in related anatomical structures, such as the ear canal, due to interconnected tissue and bone dynamics.
For example, jaw movements involved in expressions like surprise or joy can affect the mandibular condyle, leading to changes in the shape of the ear canal. Similarly, the temporalis muscle, a major muscle in the head, transmits mechanical artifacts from facial expressions toward the ear area.
An Inertial Measurement Unit (IMU) embedded in an ear-mounted device, like a wireless earbud, can be employed to capture these changes. As facial expressions lead to deformations of the ear canal, the earbud’s position and orientation are altered. The IMU detects these variations through accelerometers and gyroscopes, translating these subtle physical changes into measurable data. Each facial expression has a unique motion pattern, differentiating distinct facial expressions based on the IMU’s output signals. By employing advanced machine learning algorithms, these signals can be analyzed to identify specific facial expressions, offering a practical method to decode them via ear-mounted sensors. This approach opens new possibilities for unobtrusive facial expression recognition, leveraging the natural connection between facial musculature and ear canal dynamics.
Figure 3: Device Prototype: Composed of a pair of earbuds and XLAO nRF52840 Sense MCU.
IV Method
In Fig. 2, we present an overview of our proposed IMUFace system. First, we capture changes in facial expressions using inertial measurement units (IMUs) embedded in earphones, in conjunction with corresponding video data of the human face. The collected data undergoes preprocessing. A quick calibration process is conducted for the IMU data to ensure that the IMUs effectively capture skin movement. Next, we apply the Short-Time Fourier Transform (STFT) to extract frequency domain features from the IMU data, after which we segment this data.
To process the video data, we begin by extracting facial landmarks and normalizing them. These normalized landmarks serve as the ground truth during the training process. Next, we input the preprocessed IMU data into a transformer-based deep learning model. This model uses information from the IMU sensors and analyzes the relationships between consecutive frames to estimate the positions of facial landmarks in the current frame. Finally, we fit these landmarks to the parameters of the FLAME model (Faces Learned with an Articulated Model and Expressions) to create 3D facial reconstruction animations.
IV-A Hardware Prototype Design
We constructed an earphone-based, proof-of-concept prototype capable of capturing IMU signals generated by facial expression movements. It primarily consists of two parts: a tiny sensor board and an earphone frame. We selected the XIAO nRF52840 Sense as the sensor board, with a Bluetooth chip and a 6-axis IMU chip. The Bluetooth chip supports Bluetooth 5.0 for data transmission and directly interfaces with the IMU chip, extracting data using the I2C protocol. The sensor board measures 21.0 mm × 17.5 mm. It provides a compact and efficient solution with low power consumption, meeting our requirements perfectly.
The earphone frame we chose is a standard wireless Bluetooth earphone found in daily life, which can be any standard earphone, ensuring the broad applicability of the system. We stream the 6-axis IMU data from both earphones at a sampling rate of 30 Hz with a 16-bit resolution, ensuring the comprehensiveness of the collected muscle information. The schematic diagram of the device is shown in Fig. 3. First, we attach the battery to the back of the XIAO nRF52840 Sense, then mount it on the earphone. Actually, many modern wireless earphones/earbuds are already equipped with the IMU module inside, which could be easily adapted for deploying the IMUFace approach.
IV-B Data Processing Pipeline
IV-B1 Calibration
Since different users may have different neutral or natural head orientations, we ask users to wear the earbuds and keep their heads naturally upright before starting the video recording. We then record 4 seconds of IMU data from all channels and take the sample mean of these values, referred to as the offset. To compensate for the drift in gyroscope measurements, we adjust the gyroscope readings of each frame by the calculated offset.
IV-B2 Signal Synchronization and Filtering
To ensure synchronization between the two modalities of data streams, we record timestamps for both video data and IMU sensor data. Using the starting timestamp of the video data as a reference, we find the closest corresponding timestamp in the IMU sensor data. Since the sampling rate of the IMU sensor data fluctuates around 30 Hz, we resample the IMU sensor data to 30 Hz. Although our hardware device includes a built-in bandpass filter for reading IMU sensor data, we apply a high-pass filter with a cutoff frequency of 0.1 Hz to improve the signal-to-noise ratio.
IV-B3 Signal Normalization
We utilize the open-source tool OpenFace [36] to extract facial landmarks from facial videos. However, the positions of these landmarks can be affected by factors such as head position, pose variations, and changes in camera angle and placement. Following the approach used in the 300W dataset [37], we standardize the facial landmarks to a consistent reference frame through a series of transformations.
First, we select the nose tip as the origin and shift all landmarks accordingly:
$P_{i}^{\prime}=(x_{i}-x_{\text{nose}},y_{i}-y_{\text{nose}})$ (1)
where $P_{i}=(x_{i},y_{i})$ are the original landmark coordinates, and $(x_{\text{nose}},y_{\text{nose}})$ are the coordinates of the nose tip.
Next, by rotating the coordinates, we align the $x$-axis to the direction from the left outer eye corner to the right outer eye corner. The rotation angle $\theta$ is calculated as:
$\theta=\arctan\left(\frac{y_{\text{right}}-y_{\text{left}}}{x_{\text{right}}-x_{\text{left}}}\right)$ (2)
where $(x_{\text{left}},y_{\text{left}})$ and $(x_{\text{right}},y_{\text{right}})$ represent the coordinates of the left and right outer eye corners, respectively.
The rotated coordinates $P_{i}^{\prime\prime}$ are obtained using the following transformation:
$P_{i}^{\prime\prime}=\begin{bmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{bmatrix}\begin{bmatrix}x_{i}^{\prime}\\ y_{i}^{\prime}\end{bmatrix}$ (3)
Finally, we normalize the distance between the outer eye corners to 1, ensuring uniform scaling. The normalization factor $d$ is given by:
$d=\sqrt{(x_{\text{right}}^{\prime\prime}-x_{\text{left}}^{\prime\prime})^{2}+(y_{\text{right}}^{\prime\prime}-y_{\text{left}}^{\prime\prime})^{2}}$ (4)
The final normalized coordinates $P_{i}^{\prime\prime\prime}$ are computed as:
$P_{i}^{\prime\prime\prime}=\frac{P_{i}^{\prime\prime}}{d}$ (5)
Fig. 4 shows the facial landmark coordinates before and after alignment.
Figure 4: Illustration of facial landmark alignment.
IV-B4 Short-Time Fourier Transform
Each facial expression movement is associated with unique frequency components. Since the STFT better isolates noise components and enhances the clarity of extracted features from time-series data [38], we use STFT to convert time-domain signals into time-frequency domain spectrograms. The STFT of a signal $x(t)$ is computed as:
[…]
where $w(\tau-t)$ is a window function centered at $t$, $f$ is the frequency, and $X(t,f)$ represents the time-frequency representation of the signal. Frequency-domain features improve the training of our model. We apply STFT to each channel of the IMU data with a window length of 30, a step size of 1, and 32 frequency components. Considering the model complexity, we concatenate six spectrograms (sharing the same time axis) to form a single feature representation input. The concatenation is along the frequency axis, resulting in a final frequency-domain feature dimension of $12\times 17=204$, with a length of IMU signal length.
IV-B5 Data Segmentation
We collect IMU signal data with 12 channels and extract frequency-domain features for each channel. Both the time-domain data and the extracted frequency-domain features are used as inputs to the model. To enable the model to utilize rich temporal relationships when predicting facial landmarks, we input ten continuous frames of data, each with a size of 1 × 216.
Figure 5: Illustration of the model structure.
IV-C Model Architecture and Training Strategy
In this study, we designed a deep-learning model called IMUTwinTrans. This model, based on a ConvTransformer architecture, is capable of simultaneously leveraging the temporal and frequency features of IMU signals to establish the correspondence between IMU signals and facial landmarks, as illustrated in Fig. 5.
IV-C1 CNN-based Feature Extraction
We utilize a CNN to extract robust features from the input data frames. The architecture consists of sequentially stacked 1D convolutional layers with residual connections to enhance feature representation. The network begins with an initial convolutional block followed by three residual blocks that progressively increase the number of channels from 64 to 512 while maintaining a consistent kernel size of 3 and a stride of 1. This configuration allows the network to capture both fine-grained and hierarchical information. To mitigate overfitting, we incorporate dropout layers with a probability of 0.15 after each convolutional layer.
Additionally, each residual block includes shortcut connections to preserve gradient flow during training, improving model stability and convergence. Batch normalization is applied after each convolution to ensure faster training and better generalization. The input frames of size 10×216 are passed through this network, yielding an output of size 10×512, effectively capturing informative features for downstream tasks.
IV-C2 Transformer Structure
We perform value embedding and position encoding on the feature data derived from the CNN output (size 10×512) and feed it into a two-layer Transformer Encoder. The encoder includes a self-attention module within the multi-head attention layers, a multilayer perceptron, and normalization layers. Residual connections are incorporated at both the multi-head attention and multilayer perceptron layers to boost the convergence of the machine learning model. Regarding model parameters, we set $d_{model}$ to 512, $n_{head}$ to 4, $d_{ff}$ to 1024, and the dropout ratio to 0.1 in the encoder.
IV-C3 Loss Function
In the task of facial landmark prediction, standard MSE and MAE loss functions treat all errors equally, regardless of their magnitude. In particular, MSE heavily penalizes large errors while neglecting small ones, resulting in important but less active landmarks (such as pupils compared to lips) not receiving sufficient attention during training. To address this issue, we adopt the Wing Loss function [39], which applies a relatively smooth treatment to small errors while assigning higher weights to large errors. This helps the model predict points closer to the true values more accurately without overly penalizing the model for a few large errors. The loss for each facial landmark is defined as:
Wing $\text{Wing Loss}(x_{i})=\begin{cases}w\log(1+\frac{|x_{i}|}{\epsilon}),&\text{if}|x_{i}|<w\\ |x_{i}|-C,&\text{if}|x_{i}|\geq w\end{cases}$ (7)
where $|x_{i}|$ is the L2 distance between the ground truth and the reconstructed coordinate for the $i$ landmark. $w$ represents the threshold of the small error, which is set to 20 in our case. $\epsilon$ means the curvature in the small error range, and $C=w-wln(1+w/\epsilon)$ which links the linear part and non-linear part together.
IV-C4 Optimization
In addition, the network is trained using the Adam optimizer. The learning rate scheduler uses a Warmup + Cosine Annealing strategy. Initially, during the Warmup phase, the learning rate gradually increases to stabilize gradient updates and prevent divergence. This is followed by the Cosine Annealing phase, where the learning rate smoothly decreases following a cosine curve, allowing for more precise parameter adjustments and reducing overfitting. This strategy accelerates convergence while improving the model’s generalization.
IV-D 3D Reconstruction
To improve the system’s usability and reduce the modeling’s complexity, we seek a compact head model that can easily fit the data while preserving enough details to generate expressive facial animations.
Figure 6: Performance of continuous facial landmark tracking for each participant.
FLAME 3D Head Model. IMUFace uses the FLAME [40] model to represent the face in 3D. FLAME is a statistical head model that learns from approximately 33,000 3D scans. It combines a PCA space for identity shape, simple rotational degrees of freedom for the neck, jaw, and eyeballs, pose-dependent corrective blend shapes, and global expression blend shapes to efficiently represent complex facial geometry and expressions. Mathematically, the model is defined as:
$FLAME(\vec{\beta},\vec{\theta},\vec{\psi}):\mathbb{R}^{|\vec{\beta}|\times|\vec{\theta}|\times|\vec{\psi}|}\to\mathbb{R}^{3N}$ (8)
Optimization-based Shape Reconstruction. We employ a series of optimization steps to obtain a 3D face model that closely resembles the user’s real expressions. First, we perform camera calibration by optimizing the parameters for scale, rotation, and translation to minimize the L2 distance between the 2D projections of the 3D face model’s facial landmarks and the user’s facial landmarks. Second, we optimize the FLAME model parameters (shape, pose, and expression) to further reduce the L2 distance between the 2D projections of the facial landmarks from the 3D face model and the user’s facial landmarks. Ultimately, we can render the user’s 3D face model using the optimal FLAME model parameters.
V Performance Evaluation
To validate the effectiveness of our system, we conducted extensive and detailed user studies to collect training data and evaluate the performance of our system in facial reconstruction. We also performed a series of micro-benchmark tests using the trained model to further ensure the system’s reliability and better understand its capabilities. These tests were designed to examine the system’s performance across various scenarios.
V-A Experimental Methodology
V-A1 Data Collection
In the study, we recruited 12 participants to evaluate IMUFace’s performance. The participants included nine males and three females from the local institution, aged from 18 to 28. The experimental equipment comprises the IMUFace prototype, a Hikvision camera, and a desktop computer. Participants were asked to wear the IMUFace wireless earphones prototype, which was connected to the computer via Bluetooth.
To evaluate the performance of tracking 51 facial landmarks, we focused on seven universal facial expressions of emotion involving happiness, sadness, anger, surprise, fear, disgust, and contempt. Before the experiment, we showed the participants examples of the seven facial expressions and had them repeatedly practice making the corresponding facial expressions while wearing our implemented IMUFace prototype.
Each expression was separated by a neutral facial expression (e.g., a relaxed facial expression). To assist participants with their data collection, seven pictures of the corresponding faces were displayed on a screen for them to imitate. The pace and extent to which each expression was performed were not controlled throughout the experiments. While the users performed the facial motion, IMU data from both earphones were collected at a sampling rate of 30 Hz. To validate the accuracy of IMUFace, the built-in camera recorded videos of the user’s face on the computer at 720p resolution and 30 fps.
Each participant completed 12 sessions of the experiment. In each session, the participants were required to make seven expressions, which were prompted in a random order. Each expression was made twice. Each data collection session lasted for 60 seconds, and the participants took adequate rest between sessions. The earphone prototype was removed and remounted across sessions to validate the robustness to natural changes in sensor positions during daily usage.
V-A2 Experimental Settings
The training was performed on a desktop with an Intel i9-12700K CPU, 32 GB of RAM, and an Nvidia GTX 4090 GPU. We used the Adam optimizer with a learning rate of 0.001. To avoid overfitting issues that may happen in the training process, we added dropouts with a parameter of 0.1 following each RELU activation.
V-A3 Evaluation Metric
The IMUFace system is evaluated primarily by two metrics [19]:
1) Mean Absolute Error (MAE): the absolute difference between the predicted facial landmarks and the ground-truth facial landmarks. It is defined as: $MAE=\frac{1}{N}\times\sum\|g-r\|\times\frac{d_{real}}{d_{norm}}$, where $g$ and $r$ denote the facial landmarks extracted from the ground truth of vision and IMUFace in a normalized coordinate frame. ${d_{real}}$ denotes the actual outer corner distance in millimeters(we measured ${d_{real}}$ for each user who participated in the study). ${d_{norm}}$ denotes the outer corner distance in the normalized coordinate frame. From Section IV-B3, we know that ${d_{norm}}=1$. This provided the MAE in units of millimeters, which is the primary metric used to evaluate IMUFace.
2) Normalized Mean Error (NME): the average error between the predicted facial landmarks and the ground truth facial landmarks, represented by $NME=\frac{1}{N}\times\sum\frac{\|g-r\|}{d_{norm}}\times 100\%$, where $g$ and $r$ denote the facial landmarks extracted from the ground truth of vision and IMUFace in a normalized coordinate frame. ${d_{norm}}$ denotes the outer corner distance in the normalized coordinate frame. From Section IV-B3, we know that ${d_{norm}}=1$.
Figure 7: Domain adaptation.
TABLE I: Comparison result of different Models.
Models Average MAE Average NME
Bi-LSTM [41] 2.57 mm 3.97%
TCN [42] 2.39 mm 3.68%
IF-ConvTransformer [43] 2.27 mm 3.50%
Mamba2 [44] 2.37 mm 3.64%
IMUTwinTrans 2.21 mm 3.40%
V-B User-Dependent Evaluation
For each user’s 12 sessions data, we randomly selected five sessions for training. The rest was used for testing in a randomized cross-validation fashion. Fig. 6(a) illustrates the average MAE & NME and corresponding standard deviations for all the 51 facial landmarks of each participant. Specifically, IMUFace obtains an average of 2.21 mm MAE and 3.40% NME with average standard deviations of 0.99 mm and 0.90%. Among all participants, U4 performed the best, with an MAE of only 1.46, while U8 had the highest MAE at 2.72. Fig. 6(b) depicts the Cumulative Density Function (CDF) of the MAE errors for each individual participant as well as cross-participant cases. 60.0% of the MAEs are less than 2.21, indicating a high level of prediction accuracy. Moreover, different landmarks may exhibit varying error levels due to their differing movement characteristics. Fig. 6(c) visualizes the average MAE for all 51 major landmarks. We found that the reconstructed landmarks on the eyebrows and mouth have relatively more significant errors, while the facial landmarks around the eyes have more minor errors. This is mainly because the eyebrows and mouth change more during facial expression variations.
V-C User-Adaptive Evaluation
Due to the need for improvement in cross-user testing results, we attempted to introduce domain adaptation techniques. First, we trained a model using data from other users, then fine-tuned the model’s linear layers with 2-minute data from a specific user, and finally tested the model using the data from the same user. Fig. 7 compares the performance differences between the user-dependent and domain-adapted models. It can be observed that the performance of the domain-adapted model is significantly improved compared to the user-dependent model, with an average MAE of 2.05 mm for the specific user. This improvement may be attributed to the unique characteristics of each user’s facial muscle movements. Through fine-tuning, the model can learn specific patterns from the target user’s data, thereby enhancing its prediction performance.
V-D Micro-benchmark Tests
V-D1 Model Selection
To verify the effectiveness of our model, we compared it with several deep learning network methods commonly used for IMU data analysis to demonstrate the superiority of our model. We evaluated the performance of different networks under user-dependent experimental settings, and the results are presented in Table I.
Figure 8: Number of earbuds.
Figure 9: Training size vs error.
V-D2 Number of Earphones
To validate the system’s efficacy under single-earphone conditions, we investigated the accuracy when using only a single earphone. As shown in Fig. 8, although using both earphones can provide the best accuracy by integrating sensor data from both ears, the accuracy with individual earphones (left or right) is also very close (2.40 mm, 2.36 mm).
V-D3 Impact of Training Data Size
In practical applications, the amount of data that needs to be collected from users for training can affect user experience. Therefore, it is necessary to explore further the relationship between system performance and the required training data to minimize the burden of data collection on users. Fig. 9 shows the overall system performance when the training data size varies from 1 minute to 5 minutes for each participant. We observe that with 1 minute of training data, the accuracy already reaches 2.66 mm, and with 5 minutes of training data, the accuracy saturates at around 2.21 mm. It is shown that IMUFace can still achieve a reasonable level of accuracy even with a relatively short period of user data collection, which offers a high level of user-friendliness.
V-D4 Latency and Power Consumption
The inference time for the 51 facial landmarks was measured on a single NVIDIA GTX 4090 GPU. Our designed model requires only about 1.26 ms to reconstruct a single frame, which is sufficient for real-time applications.
In addition, we used a power monitoring device to measure the power consumption of IMUFace. Specifically, when the system is idle, that is, the microcontroller unit (MCU) is in sleep mode and does not transmit data via Bluetooth, the average power consumption of IMUFace is 50 mW (13.5 mA @ 3.70 V). When IMUFace actively collects IMU data and transmits it via Bluetooth, the total system power consumption is 58 mW (15.7 mA @ 3.70 V). This indicates that with a 100 mAh lithium polymer battery, IMUFace can continuously record data for up to 6.5 hours, meeting the needs of most applications. We examined the power consumption of recent studies on 3D facial reconstruction and summarized the overall power consumption of the system, including the use of sensing and communication power, in Table II. Our analysis shows that IMUFace consumes the least power among all sensing technologies evaluated, with energy consumption 58.0% lower than that of the next most efficient technology, Bioface-3D [19].
TABLE II: Comparison of System Power Consumption among Related Existing Studies
Studies Sensors Power Consumption
NeckFace [12] Cameras 4 W
C-Face [22] Cameras >4 W
BioFace-3D [19] Biosensors 138 mW
EarIO [20] Acoustics 153.7 mW
EyeEcho [15] Acoustics 167 mW
IMUFace IMUs 58 mW
VI Discussion
Our study aims to use conventional motion sensors, such as accelerometers and gyroscopes, to collect data. These sensors are practical and portable for use in any environment. However, the facial expression movement characteristics of users differ significantly, leading to noticeable variations in model performance across different users. Therefore, the system requires a small amount of data from the target user.
VI-A User-independent Model.
Through the experimental results, we found that the performance of the model after fine-tuning with user data is significantly better than that of the user-independent model. This is because the intensity, response, and sensitivity of facial movement data can vary significantly from person to person, leading to substantial differences between data domains. However, this may reduce usability, as new users must go through a registration phase before using the system. Our future research will explore generating large amounts of virtual IMU data using data from other modalities to train a more generalized user-independent model.
VI-B Effects of Body Movements.
In the system design, we considered the impact of initial head pose changes on the IMU signals. However, repetitive head bounces during activities such as walking can produce significant inertial signal noise [45], which can interfere with the normal operation of the system. Future research can explore the use of targeted filtering methods to minimize the noise impact introduced by such activities.
VII Conclusion
This article proposes and designs IMUFace, a system that reconstructs the 3D facial expressions of users using IMUs that are attached to (and potentially built into) wireless earphones. The system can detect the motion changes associated with facial expressions and achieve user-friendly, low-effort, continuous 3D facial reconstruction. We extracted frequency-domain features from the data and designed a lightweight deep learning model, IMUTwinTrans, based on a ConvTransformer architecture to achieve 3D facial reconstruction by leveraging both temporal and frequency features of the IMU signal. It achieves an average facial landmark tracking accuracy of 2.21 mm across 12 users. The deep learning model inference time is only 1.26 ms, satisfying real-time requirements. Our device consumes only 58 mW of power during operation and can run for 6.25 hours on a 3.7 V, 100 mAh battery. Numerous applications in the areas of affective computing, emotional well-being monitoring, facial expression recognition, AR/VR, accessibility, and user interfaces can be further explored in future work.